OpenStudy (asapt):

Find a possible solution to the equation sin (3x+13)=cos(4x)

OpenStudy (asapt):

@LegendarySadist

OpenStudy (asapt):

its part of my geometry course @LegendarySadist trig is one of the sections lol

OpenStudy (asapt):

@misssunshinexxoxo @ybarrap

OpenStudy (asapt):

nah at my schools it goes algebra 1 geometry algebra 3 and pre calc or discrete math im in geometry summer school

OpenStudy (asapt):

*algebra 2

OpenStudy (freckles):

use a co-function identity

OpenStudy (freckles):

for example, recall one of the co-function identities is: \[\cos(u)=\sin(\frac{\pi}{2}-u)\]

OpenStudy (asapt):

wow that really confused me

OpenStudy (freckles):

why does it confuse you?

OpenStudy (asapt):

because this stuff is like another language to me

OpenStudy (freckles):

but why does the identity confuse you?

OpenStudy (freckles):

is it because you never seen it?

OpenStudy (asapt):

kinda this is my first time learning this

OpenStudy (freckles):



OpenStudy (freckles):

could you find cos(theta) and sin(90-theta) using this right triangle ?

OpenStudy (freckles):